OpenStudy (anonymous):

if x+10 and y+3x are congruent, solve for x and y

OpenStudy (anonymous):

no idea, can't figure out how to solve with given only .. maybe @AkashdeepDeb can help xD

OpenStudy (akashdeepdeb):

Do you mean to say that x+10 = y+3x ?

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

x in terms of y .. and y in terms of x?

OpenStudy (anonymous):

yah i guess

OpenStudy (anonymous):

then you can solve it xD

OpenStudy (akashdeepdeb):

x+10 = y+3x 10 - y = 2x Get 1) y in terms of x and 2) x in terms of y y = 10 -2x Try to find the x part now!! :D

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ok i got y=-2x+10 already, so do i plug it in (y+3x)

OpenStudy (anonymous):

10 - y = 2x, in this equation, solve for x in terms of y.

OpenStudy (anonymous):

ok i'm confused

OpenStudy (akashdeepdeb):

Study the topic first and then answer the questions.

OpenStudy (anonymous):

i've done this kinda problem before but it always stumps me

OpenStudy (anonymous):

bye